PHONON WING SPECTRA OF PHTHALOCYANINES IN LIQUID HELIUM

\underline{B. SARTAKOV}, General Physics Institute RAS, Vavilov str. 38, 119991 Moscow, Russia; A. VILESOV, M. SLIPCHENKO, Department of Chemistry, University of Southern California, Los Angeles, CA 90089, USA; and R. LEHNIG, Institut fur Physikalische und Theoretische Chemie Universitat Regensburg, 93040 Regensburg, Germany.

LIF spectra of the (S1←S0) band of phthalocyanine and Mg-phthalocyanine in liquid^{4}He droplets were measured. Phonon wing of 0-0 line shows discret structure suggesting localization of some helium atoms in the neighborhood of the molecules. The spectral structure close to the band origin (ΔE < 3 cm^{−1}) is assigned to the collective motion of few He atoms on the molecular surface.